OpenStudy (anonymous):

your family goes to a restaurant for dinner there are 6 people in your family. some order the chicken dinner for 14.80 and some order the steak dinner for 17$ if the total bill was 91$ how many people ordered each type of dinner? solve with system of equations..

OpenStudy (anonymous):

Note that the total bill is an integer amount. If we assume that c is the number of people who order a chicken dinner, we can say that c, 2c, 3c, 4c, nor 6c do not give us integer values for the total bill. Therefore there must have been five people who ordered the chicken bill. Checking, 5(14.80) + 17 = 74 + 17 = 91.

OpenStudy (anonymous):

C + S = 6 14.8C + 17 S = $91 => 14.8C + 17 ( 6 - C ) = 91 Can you solve from here?